# A Puzzle About General Covariance and Gauge

### v2

@@ 2 What is General Covariance?
For our purposes, the conceptual core of general covariance is the requirement that the principal relationships posited by a theory must be preserved under coordinate transformations. In other words, whether some physically meaningful assertion holds true cannot depend on a choice of coordinate system; were some relationship to hold when expressed in one coordinate system, then the same relationship, expressed in different coordinates using suitably transformed representations of the objects involved in the assertion, would still hold (and in fact, express the same fact). From a modern perspective (c.f. Misner et al., 1973; Wald, 1984), the requirement that some object has the appropriate transformation properties under coordinate transformations to enter into such relationships is often replaced by the requirements that the object admit a coordinate-independent characterization and that the relationships involving such objects can be expressed using just these coordinate-independent objects.333 Our approach is especially close to Misner et al. (1973, p. 48), who identify general covariance with the requirement that “every physical quantity must be describable by a geometric object” in the sense of Nijenhuis (1952); and they attribute the first clear articulation of this view to Veblen and Whitehead (1932). (See also Misner et al., 1973, pp.302-3.) These remarks are very closely connected to our arguments here and, especially, in section 4. Coordinate transformations, then, may be re-interpreted as implementing diffeomorphisms acting on these coordindate-independent relationships, and invariance under coordinate changes can be interpreted as the requirement that physically meaningful relationships between coordinate-independent objects preserve their truth value under the action of diffeomorphisms.
@@ 4 Naturality, Geometricity, and Covariance
As we have now seen, there is a sense in which general relativity meets certain necessary conditions for general covariance, but that Yang-Mills theory does not. What is the difference? One way to understand what is going on here is to introduce the concept of a natural bundle. This is a way of capturing the idea of a “geometric object”, developed by Schouten and Haantjes (1936), Nijenhuis (1952) and others in the middle part of the century. [^7] Recall fn. 3. The basic observation is that the bundles encountered in general relativity typically are natural, whereas the ones encountered in Yang-Mills theory are not natural. [^8] Though we do not discuss them here, spinors are an example of a type of structure sometimes encountered in (extensions to) general relativity that are not natural, because to equip a manifold with a spinor structure involves an additional choice of convention (c.f. Fatibene and Francaviglia, 2003). We are grateful to Henrique Gomes to raising this point, but we will postpone further discussion of it to future work.
@@
This framework also allows us to restate the claims of section 3. Neither the principal bundles nor the associated vector bundles encountered in Yang-Mills theory are natural bundles. In fact, there are several barriers to naturality. One is that when we define the bundles used in Yang-Mills theory, it is common to specify only the fiber type, and not the global topology of the bundle. We say, for instance, that we are considering an $SU(2)$ theory, which implies the fibers are $SU(2)$ torsors, but implies nothing about the bundle topology. But given a base space and typical fiber, there are generally many principal bundles (respectively, vector bundles) available. Which bundle one chooses will determine the space of global field configurations. This failure of uniqueness, meanwhile, creates problems for defining the bundles using a functor, since a functor must assign at most one object in its codomain to each object in its domain.
Now, admittedly, the bare technical problem can be overcome, for instance by specifying that one is considering only trivial bundles. (Though to be clear, restricting to trivial bundles does not restore naturality; it simply removes one barrier.) This can always be done, but it comes at an interpretational cost, because it effectively rules out certain classes of global field configuration associated with non-trivial bundles (or, more generally, ones with different global topology than those in the codomain of the functor). More generally, if one wishes to allow that field configurations associated with different bundles over a single base space are all in some sense possible configurations of a field with Yang-Mills charge, then one cannot take those fields to be sections of a (single) natural bundle over that base space. [^11] There are deep and under-explored issues, here, about global topology of principal bundles and the physical possibility of certain global configurations of matter that we are setting aside. We are not aware of a physical situation in which the representational freedom afforded by using non-trivial principal bundles is needed. Our point is only that a generic base space and typical fiber do not even uniquely determine a principal bundle, much less a way for smooth maps on the base space to lift to the total space.
@@ 5 Gauge Naturality
These analyses bear revisiting in light of the role of principal bundles in defining gauge natural bundles. While it is true that principal bundles do not represent possible states of matter directly, the role they do play in Yang-Mills theory is nonetheless robust and important. What we have seen in this section is that the vector bundles in which matter takes its properties are, in the sense described above, determined by the principal bundles to which they are associated—in a way analogous to how natural bundles are determined by their base space. We also now see that what it means for matter valued in two different vector bundles to participate in the same Yang-Mills force on a given spacetime is for them to be images of the same principal bundle under two different gauge natural bundle functors. Likewise, the bundles associated with a given type of matter across different base spaces depends not just on the base space, but also on a choice of principal bundle over that space.
@@ 6 A Unifying Perspective
We have now seen that we can recover a sense in which Yang-Mills theory is natural, even though it is not generally covariant in the sense with which began; and we have seen some consequences for foundational discussions about the formal structure of Yang-Mills theory. But this discussion raises a further question. How should we understand the relationship between natural bundles and gauge natural bundles? There is a clear formal analogy, insofar as both are functors from some category of geometrical structures to fiber bundles associated with them. But in fact, more can be said. As we will presently argue, natural bundles can be seen as a special case of gauge natural bundles—or rather, they are gauge natural bundles that can be precomposed with functors from the category of n-manifolds to some category of principal $G$-bundles. Put another way, every natural bundle functor factors through a gauge natural bundle functor. To see this point, we will return to our motivating example of a natural bundle: the tangent bundle over $n-$manifolds. We presented the tangent bundle as something constructed from smooth manifolds, and defined the tangent bundle as a natural bundle $T:\mathcal{M}_{n}\rightarrow\mathcal{FB}$. But has often been observed, we could also think of the tangent bundle over a given manifold differently, as an associated bundle to a certain principal bundle. The relevant principal bundle in that case is the (tangent) frame bundle $GL(n,\mathbb{R})\rightarrow LM\rightarrow M$, where $GL(n,\mathbb{R})$ is the (real) general linear group in $n$ dimensions. Then, if $V$ is an $n$-dimensional vector space equipped with a representation of $GL(n,\mathbb{R})$, the associated bundle $V\rightarrow LM\times_{GL}V\rightarrow M$ is (canonically) isomorphic to the tangent bundle $TM\rightarrow M$, and sections of $V\rightarrow LM\times_{GL}V\rightarrow M$ correspond to vector fields on $M$.
@@
How can we know that every natural bundle arises in this way? It is a consequence of a classic result known as the finite order theorem (Palais and Terng, 1977). To present this, we need to introduce one more bit of machinery. [^17] Many natural bundles of interest, such as bundles of rank-$(p,q)$ tensor fields, are associated to $LM$. But other important examples of natural bundles, such as the bundle of affine connections, or more generally the bundle of $r$-jets of sections of any first-order natural bundle, are higher-order. (The bundle of affine connections, for example, is second-order.) Since these bundles become important for thinking about differential equations on a manifold, we think that this justifies the additional generality of introducing higher order frame bundles here. Let $M$ be a smooth (real) manifold of dimension $n$. An $r$-frame at $p\in M$ is an invertible $r$-jet $j^{r}_{0}f$, for some $f:\mathbb{R}^{n}\rightarrow M$ such that $f(0)=p$, where 00 denotes the zero element of $\mathbb{R}^{n}$. [^18] Recall that if $M$ and $N$ are smooth manifolds, and $f:U\rightarrow N$, $g:V\rightarrow N$ are smooth maps defined on open neighbourhoods $U$, $V$ of some $p\in M$, then $f$ and $g$ are said to be $r$-equivalent at $p$ iff they agree on all their partial derivatives up to order $r$ at $p$ (in any local coordinate charts containing $p$, $f(p)$, $g(p)$). An $r$-jet at $p$ is an equivalence class $[f]_{p}$ of smooth maps which are $r$-equivalent at $p$, and the $r$-jet at $p$ containing $f$ is denoted $j^{r}_{p}f$. For any smooth maps $f:U\subset M\rightarrow N$ and $g:V\subset N\rightarrow P$, if $f(p)=q$, then $j^{r}_{q}g\circ j^{r}_{p}f:=j^{r}_{p}(g\circ f)$. An $r$-jet $j^{r}_{p}f$, $f:U\subset M\rightarrow N$ is invertible iff there exists an $r$-jet $j^{r}_{q}g$, for some $g:V\subset N\rightarrow M$ satisfying $q=f(p)$, such that $j^{r}_{q}g\circ j^{r}_{p}f=j^{r}_{p}\mathrm{id}_{M}$ and $j^{r}_{p}f\circ j^{r}_{q}g=j^{r}_{q}\mathrm{id}_{N}$. To simplify notation in what follows, we will assume that if $j^{r}_{p}f$ is an $r$-jet, then $f$ is always a local smooth map, i.e., $f:M\rightarrow N$ denotes $f:U\subset M\rightarrow N$ for some (unspecified) open neighbourhood $U$ of $p$. We note that our definitions of $r$ jets and $r$ frames, here, makes use of $\mathbb{R}^{n}$, as opposed to a generic vector space $V$, which might be viewed as more general and/or geometric (and more consistent with the style elsewhere in the paper). But for present purposes, nothing turns on this choice, and we adopt it for simplicity and consistency with the literature. The set of all $r$-frames on $M$ $L^{r}M$ is a principal fiber bundle (the $r$-frame bundle) over $M$ with structure group $GL^{r}(n,\mathbb{R})$, where $GL^{r}(n,\mathbb{R})$ is the group of invertible $r$-jets $j^{r}_{0}f$, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ with group multiplication as composition of $r$-jets, and comes with an obvious family of projection maps $\pi^{r}_{s}$, $s\leq r$, $\pi^{r}_{s}(j^{r}_{0}f)=j^{s}_{0}f$ [^19] Our terminology follows e.g. Kolář et al. (1993), though note that $L^{r}M$ is also sometimes called the bundle of holonomic $r$-frames (to distinguish it from, e.g. the bundle of (non-holonomic) $r$-frames obtained by $r-1$ times recursively taking the bundle of first jets of sections of $LM$). It is straightforward to check that $GL^{1}(n,\mathbb{R})=GL(n,\mathbb{R})$ and $L^{1}M$ is canonically isomorphic to $LM$. Intuitively, the bundle $L^{r}M$ can be thought of as follows: whilst the fiber over $\pi(p)$ in the frame bundle $LM$ consists of all $1$-equivalence classes of smooth local homeomorphisms between $\mathbb{R}^{n}$ and $M$ at $p$, i.e. linear isomorphisms between $\mathbb{R}^{n}$ and the space of $1$-equivalence classes of curves at $p$, the fibre over $\pi(p)$ in the $L^{r}M$ consists of all $r$-equivalent smooth local homeomorphisms between $\mathbb{R}^{n}$ and $M$ at $p$.
@@
Let $M$ be a smooth manifold of dimension $n$ and let $S$ be a smooth manifold which carries a (left) $GL^{r}(n,\mathbb{R})$ action. Then the associated bundle $L^{r}M\times_{GL^{r}}S\rightarrow M$ is a natural bundle, and conversely, any natural bundle over $M$ can be constructed in this way.
@@ 7 Solder Forms
These different possible isomorphisms are not often discussed, because the standard constructions of the tangent bundle, frame bundle, and associated bundle together determine a canonical one, relative to certain choices made in those constructions. This canonical isomorphism equips the frame bundle with a solder form $\theta$, which is a linear isomorphism, at each $u\in LM$, between $T_{\pi(u)}M$ and $V$, that is equivariant with respect to the $GL(n,\mathbb{R})$ action on $LM$.
@@
As Weatherall (2016) notes, this construction is very general, in the sense that any frame bundle, including ones constructed from vector bundles that are not tangent to a manifold, comes equipped with a solder form that fixes an isomorphism between the vector bundle and associated vector bundles (of the same dimension) to its frame bundle. In more detail, let $V\rightarrow B\rightarrow M$ be any vector bundle over $M$. We can construct the frame bundle $GL(V)\rightarrow LB\xrightarrow{\wp}M$ for $B$. Since the associated bundle $LM\times_{GL}V\rightarrow M$ is isomorphic to $B$, we can equip $LB$ with an equivariant one-form that defines, at each $u\in LB$, a linear isomorphism between the fiber of $B$ at $\wp(u)$ and the fiber of $LM\times_{GL}V$ at $\wp(u)$. The construction procedures fix a preferred isomorphism, and thus a preferred solder form, in just the same way as for the tangent bundle.
@@
But not all principal bundles carry a (canonical) solder form. They do so only insofar as they are viewed as (subbundles of) the frame bundle for some particular vector bundle. And the principal bundles in Yang-Mills theory are not always thought of as frame bundles. For some authors, this fact reveals a sense in which general relativity is disanalogous to Yang-Mills theory. For example, for Anandan (1993) and Healey (2007), the lack of a solder form on principal bundles shows that Yang-Mills theory does not admit the same kind of geometrical interpretation as general relativity. Anandan even argues that the existence of the solder form “breaks” gauge invariance of the gravitational field. Healey, meanwhile, takes the solder form in general relativity to partly motivate his endorsement of a holonomy interpretation of Yang-Mills theory but not general relativity. Weatherall (2016), on the other hand, drawing on the above construction, emphasizes that every principal bundle in Yang-Mills theory can be thought of as a subbundle of a frame bundle, and in that sense does carry a solder form. For Weatherall, the important difference is that $LM$ is soldered specifically to the tangent bundle, rather than that there is a solder form at all; and he argues that this difference does not support the conclusions drawn by Healey and Anandan.
@@
Our discussion of naturality and gauge naturality in the previous sections provides a new perspective on this debate about the sense in which the solder form is a ‘real’ point of disanalogy between general relativity and Yang-Mills theory. In particular, we will show that once we have the formalism of natural and gauge natural bundles on the table, we can see that (i) solder forms are in some sense a generic feature of natural bundles, but that (ii) they are also, in a precisely analogous sense, a generic feature of gauge natural bundles. And finally, (iii) the difference between solder forms in the contexts of natural bundles and gauge natural bundles has to do with the spaces that they are soldered to, rather than the fact that there exists a solder form at all. This has two conceptual payoffs. First, it clarifies the sense in which the solder form does point to a real disanalogy between general relativity and Yang-Mills theory—namely, that the bundles one encounters in general relativity are natural rather than gauge natural. Second, it provides a deeper explication of Weatherall’s point that the important disanalogy between general relativity and Yang-Mills theory is not that there exists a solder form, but that $LM$ is soldered specifically to the tangent bundle. We will now make these claims precise.
@@
To establish point (i), we now make use of Theorem 1 along with the following fact.
@@
Fact 1 can be seen as follows. First, like $L^{1}M$, $W^{1}P$ comes equipped with a canonical form $\theta_{G}$. This is because, if $V$ is an $n$-dimensional vector space, then the associated bundle $W^{1}P\times_{W^{r}G}(V\oplus\mathfrak{g})$, where $\mathfrak{g}$ is the Lie algebra of $G$, is isomorphic to the bundle $TP/G$ (Kolář et al., 1993, p. 155). Again, there are many such isomorphisms, but the construction of the associated bundle and quotient tangent bundle determine a canonical one, relative to certain choices made in those constructions. This canonical isomorphism equips $W^{1}P$ with a canonical form $\theta_{G}$, which is a linear isomorphism, at each $u\in W^{1}P$, between $T_{\pi(u)}P$ and $V\oplus\mathfrak{g}$. Second, since $W^{1}(L^{r-1}M)$ carries the canonical form $\theta_{GL^{r-1}}$, and every local smooth map $f:\mathbb{R}^{n}\rightarrow M$ lifts to a local smooth map $j^{r}f:\mathbb{R}^{n}\times GL^{r}(n,\mathbb{R})\rightarrow L^{r}M$ in the neighbourhood of the identity element $e^{r}$ of $GL^{r}(n,\mathbb{R})$, [^20] Defined via the condition $j^{r}(j^{r}_{p}g)=j^{r}_{p}(f\circ g)$ for all $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. we can pull $\theta_{GL^{r-1}}$ back to $L^{r}M$ via the map $L^{r}M\rightarrow W^{1}(L^{r-1}M)$, $j^{r}_{0}f\rightarrow j^{1}_{(0,e^{r-1})}j^{r-1}f$ to obtain $\theta^{r}$. [^21] Note that this makes sense, since the element $j^{1}_{(0,e^{r-1})}j^{r-1}f$ depends only on $j^{r}_{0}f$. Putting Theorem 1 and Fact 1 together, we see that solder forms are a generic feature of natural bundles, in the sense that any natural bundle (excepting the 00th-order case) is associated to a bundle which carries a solder form.
[…]
Perhaps unsurprisingly, given the discussion in §6, one has the following exact analogue of Theorem 1 for gauge natural bundles, originally due to Eck (1981) and strengthened by Kolář et al. (1993)):

### v3

@@ 2 What is General Covariance?
For our purposes, the conceptual core of general covariance is the requirement that the principal relationships posited by a theory must be preserved under coordinate transformations. In other words, whether some physically meaningful assertion holds true cannot depend on a choice of coordinate system; were some relationship to hold when expressed in one coordinate system, then the same relationship, expressed in different coordinates using suitably transformed representations of the objects involved in the assertion, would still hold (and in fact, express the same fact). From a modern perspective (c.f. Misner et al., 1973; Wald, 1984), the requirement that some object has the appropriate transformation properties under coordinate transformations to enter into such relationships is often replaced by the requirements that the object admit a coordinate-independent characterization and that the relationships involving such objects can be expressed using just these coordinate-independent objects.333 Our approach is especially close to Misner et al. (1973, p. 48), who identify general covariance with the requirement that “every physical quantity must be describable by a geometric object” in the sense of Nijenhuis (1952). They attribute the first clear articulation of this view to Veblen and Whitehead (1932). (See also Misner et al., 1973, pp. 302-3.) Coordinate transformations, then, may be re-interpreted as implementing diffeomorphisms acting on these coordindate-independent relationships, and invariance under coordinate changes can be interpreted as the requirement that physically meaningful relationships between coordinate-independent objects preserve their truth value under the action of diffeomorphisms.
@@ 4 Naturality, Geometricity, and Covariance
As we have now seen, there is a sense in which general relativity meets certain necessary conditions for general covariance, but that Yang-Mills theory does not. What is the difference? One way to understand what is going on here is to introduce the concept of a natural bundle. This is a way of capturing the idea of a “geometric object”, developed by Schouten and Haantjes (1936), Nijenhuis (1952) and others in the middle part of the century. [^7] Recall fn. 3. The basic observation is that the bundles encountered in general relativity typically are natural, whereas the ones encountered in Yang-Mills theory are not natural. [^8] Though we do not discuss them here, spinors are an example of a type of structure sometimes encountered in (extensions to) general relativity that are not natural, because to equip a manifold with a spinor structure involves an additional choice of convention (c.f. Fatibene and Francaviglia, 2003). We are grateful to Henrique Gomes for raising this point, but we will postpone further discussion of it to future work.
@@
This framework also allows us to restate the claims of section 3. Neither the principal bundles nor the associated vector bundles encountered in Yang-Mills theory are natural bundles. In fact, there are several barriers to naturality. One is that when we define the bundles used in Yang-Mills theory, it is common to specify only the fiber type, and not the global topology of the bundle. We say, for instance, that we are considering an $SU(2)$ theory, which implies the fibers are $SU(2)$ torsors. But given a base space and typical fiber, there are generally many principal bundles (respectively, vector bundles) available. Which bundle one chooses will determine the space of global field configurations. This failure of uniqueness, meanwhile, creates problems for defining the bundles using a functor, since a functor must assign at most one object in its codomain to each object in its domain.
Now, admittedly, the bare technical problem can be overcome, for instance by specifying that one is considering only trivial bundles. (Though to be clear, restricting to trivial bundles does not restore naturality; it simply removes one barrier.) This can always be done, but it comes at an interpretational cost, because it effectively rules out certain classes of global field configuration associated with non-trivial bundles (or, more generally, ones with different global topology than those in the codomain of the functor). If one wishes to allow that field configurations associated with different bundles over a single base space are all in some sense possible configurations of a field with Yang-Mills charge, then one cannot take those fields to be sections of a (single) natural bundle over that base space. [^11] There are deep and under-explored issues, here, about global topology of principal bundles and the physical possibility of certain global configurations of matter. We are setting those aside. We are not aware of a physical situation in which the representational freedom afforded by using non-trivial principal bundles is needed. Our point is only that a generic base space and typical fiber do not even uniquely determine a principal bundle, much less a way for smooth maps on the base space to lift to the total space.
@@ 5 Gauge Naturality
These analyses bear revisiting in light of the role of principal bundles in defining gauge natural bundles. While it is true that principal bundles do not represent possible states of matter directly, the role they do play in Yang-Mills theory is nonetheless robust and important. What we have seen in this section is that the vector bundles in which matter takes its properties are, in the sense described above, determined by the principal bundles to which they are associated—in a way analogous to how natural bundles are determined by their base space. We also now see that what it means for matter valued in two different vector bundles to participate in the same Yang-Mills force on a given spacetime is, at least in part, for them to be images of the same principal bundle over that spacetime under two different gauge natural bundle functors. [^17] In part, because they also need to carry the “same” connection. Likewise, the bundles associated with a given type of matter across different base spaces depends not just on the base space, but also on a choice of principal bundle over that space.
@@ 6 A Unifying Perspective
We have now seen that we can recover a sense in which Yang-Mills theory is natural, even though it is not generally covariant in the sense with which we began; and we have seen some consequences for foundational discussions about the formal structure of Yang-Mills theory. But this discussion raises a further question. How should we understand the relationship between natural bundles and gauge natural bundles? There is a clear formal analogy, insofar as both are functors from some category of geometrical structures to fiber bundles associated with them. But in fact, more can be said. As we will presently argue, natural bundles can be seen as a special case of gauge natural bundles—or rather, they are gauge natural bundles that can be precomposed with functors from the category of n-manifolds to some category of principal $G$-bundles. Put another way, every natural bundle functor factors through a gauge natural bundle functor. To see this point, we will return to our motivating example of a natural bundle: the tangent bundle over $n-$manifolds, $T:\mathcal{M}_{n}\rightarrow\mathcal{FB}$. As has often been observed, the tangent bundle over a given manifold can also be thought of differently, as an associated bundle to a certain principal bundle. The relevant principal bundle in that case is the (tangent) frame bundle $GL(n,\mathbb{R})\rightarrow LM\rightarrow M$, where $GL(n,\mathbb{R})$ is the (real) general linear group in $n$ dimensions. Then, if $V$ is an $n$-dimensional vector space equipped with a representation of $GL(n,\mathbb{R})$, the associated bundle $V\rightarrow LM\times_{GL}V\rightarrow M$ is (canonically) isomorphic to the tangent bundle $TM\rightarrow M$, and sections of $V\rightarrow LM\times_{GL}V\rightarrow M$ correspond to vector fields on $M$.
@@
How can we know that every natural bundle arises in this way? It is a consequence of a classic result known as the finite order theorem (Palais and Terng, 1977). To present this result, we need to introduce another piece of machinery. [^18] Many natural bundles of interest, such as bundles of rank-$(p,q)$ tensor fields, are associated to $LM$. But other important examples of natural bundles, such as the bundle of affine connections, or more generally the bundle of $r$-jets of sections of any first-order natural bundle, are higher-order. (The bundle of affine connections, for example, is second-order.) Since these bundles become important for thinking about differential equations on a manifold, we think that this justifies the additional generality of introducing higher order frame bundles here. Let $M$ be a smooth (real) manifold of dimension $n$. An $r$-frame at $p\in M$ is an invertible $r$-jet $j^{r}_{0}f$, for some $f:\mathbb{R}^{n}\rightarrow M$ such that $f(0)=p$, where 00 denotes the zero element of $\mathbb{R}^{n}$. [^19] Recall that if $M$ and $N$ are smooth manifolds, and $f:U\rightarrow N$, $g:V\rightarrow N$ are smooth maps defined on open neighbourhoods $U$, $V$ of some $p\in M$, then $f$ and $g$ are said to be $r$-equivalent at $p$ iff they agree on all their partial derivatives up to order $r$ at $p$ (in any local coordinate charts containing $p$, $f(p)$, $g(p)$). An $r$-jet at $p$ is an equivalence class $[f]_{p}$ of smooth maps which are $r$-equivalent at $p$, and the $r$-jet at $p$ containing $f$ is denoted $j^{r}_{p}f$. For any smooth maps $f:U\subset M\rightarrow N$ and $g:V\subset N\rightarrow P$, if $f(p)=q$, then $j^{r}_{q}g\circ j^{r}_{p}f:=j^{r}_{p}(g\circ f)$. An $r$-jet $j^{r}_{p}f$, $f:U\subset M\rightarrow N$ is invertible iff there exists an $r$-jet $j^{r}_{q}g$, for some $g:V\subset N\rightarrow M$ satisfying $q=f(p)$, such that $j^{r}_{q}g\circ j^{r}_{p}f=j^{r}_{p}\mathrm{id}_{M}$ and $j^{r}_{p}f\circ j^{r}_{q}g=j^{r}_{q}\mathrm{id}_{N}$. To simplify notation in what follows, we will assume that if $j^{r}_{p}f$ is an $r$-jet, then $f$ is always a local smooth map, i.e., $f:M\rightarrow N$ denotes $f:U\subset M\rightarrow N$ for some (unspecified) open neighbourhood $U$ of $p$. We note that our definitions of $r$ jets and $r$ frames, here, makes use of $\mathbb{R}^{n}$, as opposed to a generic vector space $V$, which might be viewed as more general and/or geometric (and more consistent with the style elsewhere in the paper). But for present purposes, nothing turns on this choice, and we adopt it for simplicity and consistency with the literature. The set of all $r$-frames on $M$ $L^{r}M$ is a principal fiber bundle (the $r$-frame bundle) over $M$ with structure group $GL^{r}(n,\mathbb{R})$, where $GL^{r}(n,\mathbb{R})$ is the group of invertible $r$-jets $j^{r}_{0}f$, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ with group multiplication as composition of $r$-jets, and comes with an obvious family of projection maps $\pi^{r}_{s}$, $s\leq r$, $\pi^{r}_{s}(j^{r}_{0}f)=j^{s}_{0}f$ [^20] Our terminology follows e.g. Kolář et al. (1993), though note that $L^{r}M$ is also sometimes called the bundle of holonomic $r$-frames (to distinguish it from, e.g. the bundle of (non-holonomic) $r$-frames obtained by $r-1$ times recursively taking the bundle of first jets of sections of $LM$). It is straightforward to check that $GL^{1}(n,\mathbb{R})=GL(n,\mathbb{R})$ and $L^{1}M$ is canonically isomorphic to $LM$. Intuitively, the bundle $L^{r}M$ can be thought of as follows: whilst the fiber over $\pi(p)$ in the frame bundle $LM$ consists of all $1$-equivalence classes of smooth local homeomorphisms between $\mathbb{R}^{n}$ and $M$ at $p$, i.e. linear isomorphisms between $\mathbb{R}^{n}$ and the space of $1$-equivalence classes of curves at $p$, the fiber over $\pi(p)$ in $L^{r}M$ consists of all $r$-equivalent smooth local homeomorphisms between $\mathbb{R}^{n}$ and $M$ at $p$.
@@
[…]
@@ 7 Solder Forms
As Weatherall (2016) notes, this construction is very general, in the sense that any frame bundle, including ones constructed from vector bundles that are not tangent to a manifold, comes equipped with a solder form that fixes an isomorphism between the vector bundle and associated vector bundles (of the same dimension) to its frame bundle. In more detail, let $V\rightarrow B\rightarrow M$ be any vector bundle over $M$. We can construct the frame bundle $GL(V)\rightarrow LB\xrightarrow{\wp}M$ for $B$. Since the associated bundle $LB\times_{GL}V\rightarrow M$ is isomorphic to $B$, we can equip $LB$ with an equivariant one-form that defines, at each $u\in LB$, a linear isomorphism between the fiber of $B$ at $\wp(u)$ and the fiber of $LB\times_{GL}V$ at $\wp(u)$. The construction procedures fix a preferred isomorphism, and thus a preferred solder form, in just the same way as for the tangent bundle.
@@
But not all principal bundles carry a (canonical) solder form. They do so only insofar as they are viewed as (subbundles of) the frame bundle for some particular vector bundle. And the principal bundles in Yang-Mills theory are not always thought of as frame bundles. For some authors, this fact reveals a sense in which general relativity is disanalogous to Yang-Mills theory. For example, for Anandan (1993) and Healey (2007), the lack of a solder form on principal bundles shows that Yang-Mills theory does not admit the same kind of geometrical interpretation as general relativity. Anandan even argues that the existence of the solder form “breaks” gauge invariance for the gravitational field. Healey, meanwhile, takes the solder form in general relativity to partly motivate his endorsement of a holonomy interpretation of Yang-Mills theory but not general relativity. Weatherall (2016), on the other hand, drawing on the above construction, emphasizes that every principal bundle in Yang-Mills theory can be thought of as a subbundle of a frame bundle, and in that sense does carry a solder form. For Weatherall, the important difference is that $LM$ is soldered specifically to the tangent bundle, rather than that there is a solder form at all; and he argues that this difference does not support the conclusions drawn by Healey and Anandan.
@@
Our discussion of naturality and gauge naturality in the previous sections provides a new perspective on this debate. In particular, we will show that once we have the formalism of natural and gauge natural bundles on the table, we can see that (i) solder forms are in some sense a generic feature of natural bundles, but that (ii) they are also, in a precisely analogous sense, a generic feature of gauge natural bundles. And finally, (iii) the difference between solder forms in the contexts of natural bundles and gauge natural bundles has to do with the spaces that they are soldered to, rather than the fact that there exists a solder form at all. This has two conceptual payoffs. First, it clarifies the sense in which the solder form does point to a real disanalogy between general relativity and Yang-Mills theory—namely, that the bundles one encounters in general relativity are natural rather than gauge natural. Second, it provides a deeper explication of Weatherall’s point that the important disanalogy between general relativity and Yang-Mills theory is not that there exists a solder form, but that $LM$ is soldered specifically to the tangent bundle.
[…]
@@
We will now turn to claims (i)-(iii). To establish point (i), we now make use of Theorem 1 along with the following fact.
@@
Fact 1 can be seen as follows. First, notice that, like the frame bundles discussed in the previous section, $LM\rightarrow M$, for any principal bundle $P\rightarrow M$, $W^{1}P$ comes equipped with a canonical form $\theta_{G}$. This is because, if $V$ is an $n$-dimensional vector space, then the associated bundle $W^{1}P\times_{W^{r}G}(V\oplus\mathfrak{g})$, where $\mathfrak{g}$ is the Lie algebra of $G$, is isomorphic to the bundle $TP/G$ (Kolář et al., 1993, p. 155). Again, there are many such isomorphisms, but the construction of the associated bundle and quotient tangent bundle determine a canonical one, relative to certain choices made in those constructions. This canonical isomorphism equips $W^{1}P$ with a canonical form $\theta_{G}$, which is a linear isomorphism, at each $u\in W^{1}P$, between $T_{\pi(u)}P$ and $V\oplus\mathfrak{g}$. Now observe that for any $r\geq 1$, $L^{r}M\rightarrow M$ is a principal bundle, and so we can apply the principal prolongation construction just described. It follows that $W^{1}(L^{r-1}M)$ carries the canonical form $\theta_{GL^{r-1}}$. Since every local smooth map $f:\mathbb{R}^{n}\rightarrow M$ lifts to a local smooth map $j^{r}f:\mathbb{R}^{n}\times GL^{r}(n,\mathbb{R})\rightarrow L^{r}M$ in the neighbourhood of the identity element $e^{r}$ of $GL^{r}(n,\mathbb{R})$, [^23] Defined via the condition $j^{r}(j^{r}_{p}g)=j^{r}_{p}(f\circ g)$ for all $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. we can pull $\theta_{GL^{r-1}}$ back to $L^{r}M$ via the map $L^{r}M\rightarrow W^{1}(L^{r-1}M)$, $j^{r}_{0}f\rightarrow j^{1}_{(0,e^{r-1})}j^{r-1}f$ to obtain $\theta^{r}$. [^24] Note that this makes sense, since the element $j^{1}_{(0,e^{r-1})}j^{r-1}f$ depends only on $j^{r}_{0}f$. Putting Theorem 1 and Fact 1 together, we see that solder forms are a generic feature of natural bundles, in the sense that any natural bundle is associated to a bundle which carries a solder form. [^25] To be slightly more precise, we have established this claim for natural bundles of order $\geq 0$. The 0th order case consists of natural bundles that are not gauge natural over the frame bundle, because they are just bundles of scalar fields.
We now move on to establish point (ii). The key point here is that there is a version of Theorem 1 for gauge natural bundles as well, originally due to Eck (1981) and strengthened by Kolář et al. (1993)):